find 1/2% of 1000
Im not sure how to solve this problem.. do I find 5o% of 1000 since 1/2=50
please help

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

1% of 1000 = 10
1/2 of that = 5

Answer by richwmiller(17219)

.010=1/100=1%
you want 1/2 of that
.005*1000
(5/1000)1000=5
It is easy to get lost in the decimal points and zeroes